# Rourkela - 769008, India

### v1

@@ 3.1.1 Massless particles
We now compute the Lyapunov exponent’s thermal profile associated with massless or massive particles moving in the dyonic black hole background. For the massless particle, this can be done using Eqs. (13) and (23). The Lyapunov exponent as a function of dyonic black holes’ temperature and horizon radius is shown in Fig. 4. Here, the same colour coding and parameter values as in Fig. 2 are used. When the magnetic charge $q_{m}$ is less than the critical value $q_{mc}$, the Lyapunov exponent exhibits multi-valuedness in some temperature range. As we increase $T$ from 00 to $T_{1}$, $\lambda$ decreases gradually, as shown by the blue curve in Fig. 4(a). At $T_{1}$, the Lyapunov exponent becomes multi-valued. In particular, in the temperature range $T_{1}\leq T\leq T_{2}$, there are two profiles for which the Lyapunov exponent decreases with temperature and one for which it increases with temperature. The former two profiles appear in the small and large black hole phases, whereas the latter profile appears in the intermediate phase. The Lyapunov exponents in the small, intermediate, and large black hole phases are represented by blue, red, and green lines, respectively, in Fig. 4(a). The multi-valued nature of $\lambda$ corresponds to the swallow tail in the $T-G$ diagram. For convenience, we have also shown the small/large black hole transition temperature $T_{p}$ in Fig. 4. The Lyapunov exponent becomes single-valued for $T>T_{2}$ when $q<q_{mc}$. This behaviour is again similar to the free energy behaviour. For large temperatures, the Lyapunov exponent approaches a constant value.
@@
For $q>q_{mc}$, when there is no small/large black hole phase transition, the Lyapunov exponent remains single-valued for all temperatures. In particular, it gradually decreases and then attains a constant value as we increase the temperature. This is shown in Fig. 4(b). The Lyapunov exponent’s overall behaviour as a temperature function for different magnetic charge values $q_{m}$ is shown in Fig. 4(c). Here, we can see that, just like for the free energy, the thermal profile of the Lyapunov exponent becomes multi-valued for only those values of $q_{m}$ which are less than $q_{mc}$. Importantly, this behaviour of the Lyapunov exponent, specifically the transition from multi-valuedness to single-valuedness, correctly pinpoints the second-order critical point $q_{mc}$, i.e., the critical point is reflected in the thermal behaviour of the Lyapunov exponent. This suggests that the information about the thermodynamic phase structure and transition of dyonic black holes, to some extent, are encoded in the Lyapunov exponent. For completion, in Fig. 4(d), we have shown the plot of the Lyapunov exponent as a function of horizon radius. Here, the black dot points correspond to the extremal horizon radius. We observe that the maximum value of $\lambda$ decreases as $q_{m}$ increases at the extremal $r_{h}$. This is true for all $\phi_{e}$. Moreover, for large $r_{h}$, the $\lambda$ values for different $q_{m}$ attain a constant value, which is unity for the unit AdS radius $l$ considered here. This saturation value varies as we vary the AdS-radius, but it is always unity for a dimensionless model, i.e., by scaling various variables by $l$. This is an interesting result, which is also true in other gravity models in the following sections, implying some universality in $\lambda$.
@@ 3.1.2 Massive particles
3.2 Bardeen Black Holes
Bardeen black holes are the first examples of a regular black hole satisfying the weak energy conditions bardeen1968non . They have a distinct “regular” centre, meaning they do not contain a singularity at their core, unlike the more well-known Schwarzschild and Kerr black holes. This is because matter cannot collapse to a point as a zone of repulsive gravity surrounds Bardeen black holes. The Bardeen black holes are usually obtained from the Einstein equation with nonlinear electrodynamics sources Ayon-Beato:2000mjt . The various ways to construct such regular black hole solutions can be found in Hayward:2005gi ; Bronnikov:2000vy ; Berej:2006cc ; Dymnikova:2004zc ; Ayon-Beato:1998hmi ; Bronnikov:2005gm . Although no astrophysical Bardeen black holes have been discovered to date, theoretical physicists still find them an important subject of study as they provide a unique viewpoint on the nature of black holes and the underlying laws of physics tzikas2019bardeen ; singh2020thermodynamics ; guo2020joule .
Here, we focus on the Bardeen black hole obtained in Ayon-Beato:2000mjt from a particular nonlinear electromagnetic source. The line element of this Bardeen-AdS black hole is given by
@@ 3.2 Bardeen Black Holes
where $\mathcal{T}=T/(1-\Pi)$, $\Phi_{g}=\phi_{g}/(1-\Pi)$, $\Pi=1-r_{h}^{3}(g^{2}+r_{h}^{2})^{-3/2}$, and $\Phi_{g}$ is the potential associated with the charge $g$. Notice that this first-law expression differs from the usual first-law differential expression. The difference mainly arises due to the nonlinear nature of the gauge field sources Zhang:2016ilt ; Guo:2021wcf . [^3] The standard version of the first law can also be satisfied in the Bardeen black hole. This requires logarithmic corrections to the black hole entropy Tzikas:2018cvs ; li2019thermodynamics . Unfortunately, this looks unphysical, as the nonlinear gauge field sources do not give any additional contribution to the Wald entropy. Therefore, one also expects the black hole entropy to be given by the horizon area in the Bardeen black hole.
@@
By further increasing the value of $\tilde{g}$, there appears a critical value $\tilde{g}_{c}$ at which the three black hole phases merge to form a single black hole phase, i.e., the size of the swallow-tail decreases with $\tilde{g}$ and completely vanishes at $\tilde{g}_{c}$. Therefore, the first-order small/large transition line stops at $\tilde{g}_{c}$. Analogously to the dyonic black hole case, this $\tilde{g}^{c}$ defines a second-order critical point. This implies that the thermodynamics of the Bardeen-AdS black hole is heavily affected by the magnetic charge $\tilde{g}$. Also, this behaviour is again completely analogous to the van der Waals-type phase transition of the liquid/gas system. The magnitude of the critical point $\tilde{g}_{c}$ can be further determined from the condition of the inflexion point
$\frac{\partial\tilde{T}}{\partial\tilde{r_{h}}}=0,~{}~{}\frac{\partial^{2}\tilde{T}}{\partial\tilde{r_{h}}^{2}}=0\,,$ (45)
@@
Figure 9: The 3D plot of the effective potential $V_{\text{eff}}(r)$ as a function of horizon radius $\tilde{r_{h}}$ and orbit radius $r$ of the massive particle. Here, $L=20$ and $\tilde{g}=0.05$ are used. The red and green curves projected below correspond to the unstable and stable equilibria of the time-like circular geodesics, respectively.
@@
where $L$ and $r$ are the angular momentum of the particle and the radius of the particle’s orbit respectively. We can see that $V_{\text{eff}}(r)$ is a function of $r$ and $r_{h}$, and its 3D plot is shown in Fig. 9 where we have fixed $\tilde{g}=0.05$ and $L=20$. The curve with the extrema is projected onto the $V_{\text{eff}}(r)=-1500$ plane, where the red part indicates the unstable maxima and the green part indicates the stable minima. Beyond a certain value of $r_{h}=0.884$, there are no extrema, marked by the black line on the curve. We are mainly interested in the red part of the curve, and the analysis is done in the next subsections.
3.2.1 Massless particles
@@
where $k=\tilde{r}_{h}^{2}+1$, $m=\sqrt{\tilde{g}^{2}+\tilde{r}^{2}}$ and $n=\sqrt{\tilde{g}^{2}+\tilde{r}_{h}^{2}}$, which suggests that the Lyapunov exponent depends non-trivially on parameters $\tilde{g}$ and $\tilde{r}_{h}$ of the Bardeen black hole. Our numerical results for the Lyapunov exponent are shown in Fig. 10. We again observe a distinct behaviour of $\lambda(\tilde{T})$ depending on the magnitude of $\tilde{g}$. In particular, when $\tilde{g}$ is less than the critical value $\tilde{g_{c}}$, $\lambda$ is single-valued for $\tilde{T}\in(0,\tilde{T}_{1})$, becomes multi-valued for $\tilde{T}\in(\tilde{T}_{1},\tilde{T}_{2})$, and again becomes single-valued for $\tilde{T}>\tilde{T}_{2}$. The small/large transition temperature $\tilde{T}_{p}$ is between $\tilde{T}_{1}$ and $\tilde{T}_{2}$ and is indicated by the dashed black line. Therefore, the Lyapunov exponent is single-valued in the thermodynamically favoured small and large black hole phases and exhibits multi-valuedness at the transition temperature. In contrast, when $\tilde{g}$ exceeds the critical value, $\lambda$ is single-valued, showing that only one black hole solution exists without any phase transition. In this case, the Lyapunov exponent first increases and gradually decreases to approach a constant value finally. In Fig. 10(c), we have shown the effects of the control parameter $\tilde{g}$ on the Lyapunov exponent. As we continue to increase the value of $\tilde{g}$ from $0.04$ to $0.14$, we observe that the multi-valued nature of the Lyapunov exponent gradually decreases and disappears completely beyond the critical value, which shows the typical nature of a first-order phase transition. These observations give weight to the idea that the Lyapunov exponents can also be used to study the black hole phase transitions.
@@
Figure 10: Lyapunov exponent $\lambda$ of the massless particle as a function of temperature $T$ and horizon radius $r_{h}$ for the Bardeen black hole.
@@
Figure 12: Lyapunov exponent $\lambda$ of the massless particle as a function of temperature $T$ and Horizon radius $r_{h}$ for the Bardeen black hole.
3.2.2 Massive particles
We can similarly analyse the thermal structure of the Lyapunov exponent of the massive particle in the Bardeen black hole background. The form of the corresponding effective potential also suggests that the massive Lyapunov exponent will again depend non-trivially on $\tilde{g}$ and $\tilde{r}_{h}$. Our numerical results are shown in Fig. 12. We again observe that when the magnetic charge value $\tilde{g}$ is less than the critical value $\tilde{g_{c}}$, $\lambda$ is a multi-valued function of $\tilde{T}$ in some temperature range, suggesting multiple black hole branches. On the contrary, when $\tilde{g}$ exceeds the critical value, $\lambda$ is single-valued, showing that only one black hole solution exists. In Fig. 12(c), we have demonstrated the effect of $\tilde{g}$ on the Lyapunov exponent. As we continue to increase the value of $\tilde{g}$ from $0.04$ to $0.14$, the multi-valued nature of the Lyapunov exponent gradually changes, and it disappears completely beyond the critical value $\tilde{g}_{c}$. These observations are again similar to our previous results and suggest encoding of the phase transition critical points in the massive Lyapunov exponent.
@@ 3.2.2 Massive particles
3.3 Gauss-Bonnet Black Holes
@@
Gauss-Bonnet black holes are distinct black holes that can manifest in spacetime with more than four dimensions, owing to the presence of curvature terms that go beyond the standard Einstein-Hilbert action. The Gauss-Bonnet invariant, a fourth-order polynomial derived from the Riemann curvature tensor, gives rise to these curvature terms. The Gauss-Bonnet black holes are the simplest prototype of higher-order curvature gravity theories that arise naturally in quantum gravity theories such as string theory and are expected to play an important role in quantum gravity Boulware:1985wk ; Deser:2002jk ; Gross:1986mw . In addition to the typical features of black holes, such as a singularity and an event horizon, Gauss-Bonnet black holes exhibit attractive geometric qualities. For example, the Gauss-Bonnet theorem, which links the total curvature of a closed surface to its topology, has been extended to Gauss-Bonnet black holes. Specifically, the topology of the event horizon determines the Gauss-Bonnet coupling constant, which measures the strength of the curvature terms. These black holes have been the subject of extensive research in recent years, both for their mathematical characteristics and potential relevance to astrophysics Cai:2001dz ; Cai:2013qga ; Pani:2009wy ; Clunan:2004tb ; Wei:2014hba ; Zhou:2020vzf .
@@
The Gauss-Bonnet black holes exist only in spacetime dimensions greater than four, i.e., the Gauss-Bonnet term becomes topological (or to a total derivative term) in four dimensions. It does not contribute to the field equations. However, in dimensions greater than or equal to five, the Gauss-Bonnet term does contribute to the Einstein equation and leads to interesting solutions and properties. For instance, the Gauss-Bonnet black hole solutions not only come with different horizon topologies, such as the planar, spherical, or hyperbolic topologies, but their thermodynamic phase structure also becomes much more interesting. In particular, the spherical charged Gauss-Bonnet black hole exhibits the Hawking/Page and small/large black hole phase transitions, and the associated critical points depend on the dimensions as well as the Gauss-Bonnet coupling constant Cai:2001dz ; Cai:2013qga . Similarly, the extended phase space thermodynamic of the Gauss-Bonnet black hole exhibits $P-V$ criticality and small/large black hole phase transitions in five dimensions that, depending on the magnitude of the Gauss-Bonnet coupling constant and charge, may change in higher dimensions Cai:2013qga .
@@
It is instructive to analyse the interplay of phase transitions and Lyapunov exponent in higher derivative Gauss-Bonnet black holes. This interplay was discussed in lyu2023probing for the standard thermodynamics. Here, we concentrate on this interplay in the extended-phase space thermodynamics. We will also comment on this interplay in the standard thermodynamics, thereby complementing the work of lyu2023probing and making nontrivial new observations not discussed there. For the extended phase thermodynamics of Gauss-Bonnet black holes, we follow Cai:2013qga ; Wei:2014hba . We will work in five dimensions for simplicity as the discussion can be easily generalised for higher dimensions.
@@
It is worth pointing out that the black hole mass $M$ acts as the enthalpy $H$ rather than the internal energy of the gravitational system in the extended phase thermodynamics Kastor:2009wy . Similarly, the expressions for temperature ($T$), entropy ($S$), volume ($V$), and electric potential ($\Phi$) are given by  [^4] We have set the five-dimensional Newton’s constant to one in these expressions.
@@
Figure 16: The 3D plot of the effective potential $V_{\text{eff}}(r)$ as a function of horizon radius $r_{h}$ and orbit radius $r$ for the massive particle when $L=20$, $P=0.1$, $\alpha=0.01$ and $Q=0.2$. The pink and green curves projected in the lower part correspond to the unstable and stable equilibria of the time-like circular geodesics.

### v2

@@ III.1.1 Massless particles
We now compute the Lyapunov exponent’s thermal profile associated with massless or massive particles moving in the dyonic black hole background. For the massless particle, this can be done using Eqs. (13) and (23). The Lyapunov exponent as a function of dyonic black holes’ temperature and horizon radius is shown in Fig. 4. Here, the same colour coding and parameter values as in Fig. 2 are used. When the magnetic charge $q_{m}$ is less than the critical value $q_{mc}$, the Lyapunov exponent exhibits multi-valuedness in some temperature range. As we increase $T$ from 00 to $T_{1}$, $\lambda$ decreases gradually, as shown by the blue curve in Fig. 4(a). At $T_{1}$, the Lyapunov exponent becomes multi-valued. In particular, in the temperature range $T_{1}\leq T\leq T_{2}$, there are two profiles: one for which the Lyapunov exponent decreases with temperature and one for which it increases with temperature. The former two profiles appear in the small and large black hole phases, whereas the latter profile appears in the intermediate phase. The Lyapunov exponents in the small, intermediate, and large black hole phases are represented by blue, red, and green lines, respectively, in Fig. 4(a). The multi-valued nature of $\lambda$ corresponds to the swallow tail in the $G-T$ diagram. For convenience, we have also shown the small/large black hole transition temperature $T_{p}$ in Fig. 4. The Lyapunov exponent becomes single-valued for $T>T_{2}$ when $q<q_{mc}$. This behaviour is again similar to the free energy behaviour. For large temperatures, the Lyapunov exponent approaches a constant value.
@@
For $q>q_{mc}$, when there is no small/large black hole phase transition, the Lyapunov exponent remains single-valued for all temperatures. In particular, it gradually decreases and then attains a constant value as we increase the temperature. This is shown in Fig. 4(b). The overall behaviour of the Lyapunov exponent as a function of temperature for different magnetic charge values $q_{m}$ is shown in Fig. 4(c). Here, we can see that, just like for the free energy, the thermal profile of the Lyapunov exponent becomes multi-valued for only those values of $q_{m}$ which are less than $q_{mc}$. Importantly, this behaviour of the Lyapunov exponent, specifically the transition from multi-valuedness to single-valuedness, correctly pinpoints the second-order critical point $q_{mc}$, i.e., the critical point is reflected in the thermal behaviour of the Lyapunov exponent. This suggests that the information about the thermodynamic phase structure and transition of dyonic black holes, to some extent, are encoded in the Lyapunov exponent. For completion, in Fig. 4(d), we have shown the plot of the Lyapunov exponent as a function of horizon radius. Here, the black dot points correspond to the extremal horizon radius. We observe that the maximum value of $\lambda$ decreases as $q_{m}$ increases at the extremal $r_{h}$. This is true for all $\phi_{e}$. Moreover, for large $r_{h}$, the $\lambda$ values for different $q_{m}$ attain a constant value, which is unity for the unit AdS radius $l$ considered here. This saturation value varies as we vary the AdS-radius, but it is always unity for a dimensionless model, i.e., by scaling various variables by $l$. This is an interesting result, which is also true in other gravity models in the following sections, implying some universality in $\lambda$.
@@ III.1.2 Massive particles
III.2 Bardeen Black Holes
Bardeen black holes are the first examples of a regular black hole satisfying the weak energy conditions [108]. They have a distinct “regular” centre, meaning they do not contain a singularity at their core, unlike the more well-known Schwarzschild and Kerr black holes. This is because matter cannot collapse to a point as a zone of repulsive gravity surrounds Bardeen black holes. The Bardeen black holes are usually obtained from the Einstein equation with nonlinear electrodynamics sources [89]. [^3] Recently, it has been suggested that an infinite tower of higher-curvature corrections can also give rise to regular black hole solutions [109]. The various ways to construct such regular black hole solutions can be found in [110, 111, 112, 113, 114, 115]. Although no astrophysical Bardeen black holes have been discovered to date, theoretical physicists still find them an important subject of study as they provide a unique viewpoint on the nature of black holes and the underlying laws of physics [116, 117, 118].
Here, we focus on the Bardeen black hole obtained in [89] from a particular nonlinear electromagnetic source. The line element of this Bardeen-AdS black hole is given by
@@ III.2 Bardeen Black Holes
where $\mathcal{T}=T/(1-\Pi)$, $\Phi_{g}=\phi_{g}/(1-\Pi)$, $\Pi=1-r_{h}^{3}(g^{2}+r_{h}^{2})^{-3/2}$, and $\Phi_{g}$ is the potential associated with the charge $g$. Notice that this first-law expression differs from the usual first-law differential expression. The difference mainly arises due to the nonlinear nature of the gauge field sources [119, 120]. [^4] The standard version of the first law can also be satisfied in the Bardeen black hole. This requires logarithmic corrections to the black hole entropy [121, 122]. Unfortunately, this looks unphysical, as the nonlinear gauge field sources do not give any additional contribution to the Wald entropy. Therefore, one also expects the black hole entropy to be given by the horizon area in the Bardeen black hole.
@@
By further increasing the value of $\tilde{g}$, there appears a critical value $\tilde{g}_{c}$ at which the three black hole phases merge to form a single black hole phase, i.e., the size of the swallow-tail decreases with $\tilde{g}$ and completely vanishes at $\tilde{g}_{c}$. Therefore, the first-order small/large transition line stops at $\tilde{g}_{c}$. Analogously to the dyonic black hole case, this $\tilde{g}_{c}$ defines a second-order critical point. This implies that the thermodynamics of the Bardeen-AdS black hole is heavily affected by the magnetic charge $\tilde{g}$. Also, this behaviour is again completely analogous to the van der Waals-type phase transition of the liquid/gas system. The magnitude of the critical point $\tilde{g}_{c}$ can be further determined from the condition of the inflexion point
$\frac{\partial\tilde{\mathcal{T}}}{\partial\tilde{r_{h}}}=0,~{}~{}\frac{\partial^{2}\tilde{\mathcal{T}}}{\partial\tilde{r_{h}}^{2}}=0\,,$ (45)
@@
Figure 9: The 3D plot of the effective potential $V_{\text{eff}}(r)$ as a function of horizon radius $\tilde{r_{h}}$ and orbit radius $r$ of the massive particle. Here, $L=20$ and $\tilde{g}=0.05$ are used. The red and green curves projected below correspond to the unstable and stable equilibria of the timelike circular geodesics, respectively.
@@
where $L$ and $r$ are the angular momentum of the particle and the radius of the particle’s orbit respectively. We can see that $V_{\text{eff}}(r)$ is a function of $r$ and $r_{h}$, and its 3D plot is shown in Fig. 9 where we have fixed $\tilde{g}=0.05$ and $L=20$. The curve with the extrema is projected onto the $V_{\text{eff}}(r)=-1500$ plane, where the red part indicates the unstable maxima and the green part indicates the stable minima. Beyond a certain value of $r_{h}=0.884$, there are no extrema (marked by the black line on the curve). We are mainly interested in the red part of the curve, and the analysis is done in the next subsections.
III.2.1 Massless particles
@@
where $k=\tilde{r}_{h}^{2}+1$, $m=\sqrt{\tilde{g}^{2}+\tilde{r}^{2}}$ and $n=\sqrt{\tilde{g}^{2}+\tilde{r}_{h}^{2}}$, which suggests that the Lyapunov exponent depends non-trivially on parameters $\tilde{g}$ and $\tilde{r}_{h}$ of the Bardeen black hole. Our numerical results for the Lyapunov exponent are shown in Fig. 10.
@@
Figure 10: Lyapunov exponent $\lambda$ of the massless particle as a function of temperature $\tilde{\mathcal{T}}$ and horizon radius $r_{h}$ for the Bardeen black hole.
@@
We again observe a distinct behaviour of $\lambda(\tilde{\mathcal{T}})$ depending on the magnitude of $\tilde{g}$. In particular, when $\tilde{g}$ is less than the critical value $\tilde{g_{c}}$, $\lambda$ is single-valued for $\tilde{\mathcal{T}}\in(0,\tilde{\mathcal{T}}_{1})$, becomes multi-valued for $\tilde{\mathcal{T}}\in(\tilde{\mathcal{T}}_{1},\tilde{\mathcal{T}}_{2})$, and again becomes single-valued for $\tilde{\mathcal{T}}>\tilde{\mathcal{T}}_{2}$. The small/large transition temperature $\tilde{\mathcal{T}}_{p}$ is between $\tilde{\mathcal{T}}_{1}$ and $\tilde{\mathcal{T}}_{2}$ and is indicated by the dashed black line. Therefore, the Lyapunov exponent is single-valued in the thermodynamically favoured small and large black hole phases and exhibits multi-valuedness at the transition temperature. In contrast, when $\tilde{g}$ exceeds the critical value, $\lambda$ is single-valued, showing that only one black hole solution exists without any phase transition. In this case, the Lyapunov exponent first increases and gradually decreases to finally approach a constant value. In Fig. 10(c), we have shown the effects of the control parameter $\tilde{g}$ on the Lyapunov exponent. As we continue to increase the value of $\tilde{g}$ from $0.04$ to $0.14$, we observe that the multi-valued nature of the Lyapunov exponent gradually decreases and disappears completely beyond the critical value, which shows the typical nature of a first-order phase transition. These observations give weight to the idea that the Lyapunov exponents can also be used to study the black hole phase transitions.
@@ III.2.2 Massive particles
Figure 12: Lyapunov exponent $\lambda$ of the massless particle as a function of temperature $\tilde{\mathcal{T}}$ and Horizon radius $r_{h}$ for the Bardeen black hole.
@@
We can similarly analyse the thermal structure of the Lyapunov exponent of the massive particle in the Bardeen black hole background. The form of the corresponding effective potential also suggests that the massive Lyapunov exponent will again depend non-trivially on $\tilde{g}$ and $\tilde{r}_{h}$. Our numerical results are shown in Fig. 12. We again observe that when the magnetic charge value $\tilde{g}$ is less than the critical value $\tilde{g_{c}}$, $\lambda$ is a multi-valued function of $\tilde{\mathcal{T}}$ in some temperature range, suggesting multiple black hole branches. On the contrary, when $\tilde{g}$ exceeds the critical value, $\lambda$ is single-valued, showing that only one black hole solution exists. In Fig. 12(c), we have demonstrated the effect of $\tilde{g}$ on the Lyapunov exponent. As we continue to increase the value of $\tilde{g}$ from $0.04$ to $0.14$, the multi-valued nature of the Lyapunov exponent gradually changes, and it disappears entirely beyond the critical value $\tilde{g}_{c}$. These observations are again similar to our previous results and suggest encoding of the phase transition critical points in the massive Lyapunov exponent.
@@
III.3 Gauss-Bonnet Black Holes
@@
Gauss-Bonnet black holes are distinct black holes that can manifest in spacetime with more than four dimensions, owing to the presence of curvature terms that go beyond the standard Einstein-Hilbert action. The Gauss-Bonnet invariant, a fourth-order polynomial derived from the Riemann curvature tensor, gives rise to these curvature terms. The Gauss-Bonnet black holes are the simplest prototype of higher-order curvature gravity theories that arise naturally in quantum gravity theories such as string theory and are expected to play an important role in quantum gravity [90, 91, 92]. In addition to the typical features of black holes, such as a singularity and an event horizon, Gauss-Bonnet black holes exhibit attractive geometric qualities. For example, the Gauss-Bonnet theorem, which links the total curvature of a closed surface to its topology, has been extended to Gauss-Bonnet black holes. Specifically, the topology of the event horizon determines the Gauss-Bonnet coupling constant, which measures the strength of the curvature terms. These black holes have been the subject of extensive research in recent years, both for their mathematical characteristics and potential relevance to astrophysics [123, 37, 124, 125, 126, 127].
@@
The Gauss-Bonnet black holes exist only in spacetime dimensions greater than four, i.e., the Gauss-Bonnet term becomes topological (or to a total derivative term) in four dimensions. It does not contribute to the field equations. However, in dimensions greater than or equal to five, the Gauss-Bonnet term does contribute to the Einstein equation and leads to interesting solutions and properties. For instance, the Gauss-Bonnet black hole solutions not only come with different horizon topologies, such as the planar, spherical, or hyperbolic topologies, but their thermodynamic phase structure also becomes much more interesting. In particular, the spherical charged Gauss-Bonnet black hole exhibits the Hawking/Page and small/large black hole phase transitions, and the associated critical points depend on the dimensions as well as the Gauss-Bonnet coupling constant [123, 37]. Similarly, the extended phase space thermodynamic of the Gauss-Bonnet black hole exhibits $P-V$ criticality and small/large black hole phase transitions in five dimensions that, depending on the magnitude of the Gauss-Bonnet coupling constant and charge, may change in higher dimensions [37].
@@
It is instructive to analyse the interplay of phase transitions and Lyapunov exponent in higher derivative Gauss-Bonnet black holes. This interplay was discussed in [84] for the standard thermodynamics. Here, we concentrate on this interplay in the extended-phase space thermodynamics. We will also comment on this interplay in the standard thermodynamics, thereby complementing the work of [84] and making nontrivial new observations not discussed there. For the extended phase thermodynamics of Gauss-Bonnet black holes, we follow [37, 126]. We will work in five dimensions for simplicity as the discussion can be easily generalised for higher dimensions.
@@
It is worth pointing out that the black hole mass $M$ acts as the enthalpy $H$ rather than the internal energy of the gravitational system in the extended phase thermodynamics [57]. Similarly, the expressions for temperature ($T$), entropy ($S$), volume ($V$), and electric potential ($\Phi$) are given by  [^5] We have set the five-dimensional Newton’s constant to one in these expressions.
@@
Figure 16: The 3D plot of the effective potential $V_{\text{eff}}(r)$ as a function of horizon radius $r_{h}$ and orbit radius $r$ for the massive particle when $L=20$, $P=0.1$, $\alpha=0.01$ and $Q=0.2$. The pink and green curves projected in the lower part correspond to the unstable and stable equilibria of the timelike circular geodesics.